OpenStudy (anonymous):

Write a polynomial function in standard form with zeros at 5, –4, and 1.

jimthompson5910 (jim_thompson5910):

x = 5, x = -4 or x = 1 x-5=0, x+4=0 or x-1=0 (x-5)(x+4)(x-1)=0 (x-5)(x^2-x+4x-4)=0 (x-5)(x^2+3x-4)=0 x(x^2+3x-4)-5(x^2+3x-4)=0 x^3+3x^2-4x-5x^2-15x+20=0 x^3-2x^2-19x+20=0 So the polynomial in standard form that has zeros 5, -4, and 1 is x^3-2x^2-19x+20